if a body is projected with a velocity 20 root 2 m/s making an angle 45 degree with the horizontal. Then initial velocity of the particle in vector form....

Grade:12th pass

2 Answers

Nishant Vora IIT Patna
7 years ago
20 root2 can be resolved in hori and vertical direction as 20 root2 cos45 and 20 root2 sin45.

Rahul Verma
7 years ago
20\sqrt{2} can be resolved in two components  20\sqrt{2}sin45  in vertical directon and 20\sqrt{2}cos45 in horizontal direction
so in vertical direction component =20\sqrt{2}\times \frac{1}{\sqrt{2}}=20 which can be written in vector as 20j
in horizontal direction component =20\sqrt{2 }\times \frac{1}{\sqrt{2}}=20which can be written in vector as 20i
adding both the vectors we get 20i + 20j